erin and susan both took their dogs for a walk. together they walked 15 miles. if erin walked3 milesless than susan, how far did susan walk with her dog?
Answer by orca(409)

Let x be the distance Susan walked, then the distance Erin walked is x - 3.
As the total distance walked is 15, we have:
x + x - 3 = 15
Solving for x, we obtain
2x - 3 = 15
2x = 18
x = 9
So Susan walked 9 miles, and Erin walked x - 3 = 9 - 3 = 6 miles.